### v1

@@ 9.3.1 Time-series evaluation measures
With a few exceptions, dynamical systems-based benchmarks typically output a scalar value, rather than a multidimensional vector value, each timestep. Evaluation is then typically performed using one of Mean Squared Error (MSE), Root Mean Squared Error (RMSE), Normalised Mean Squared Error (NMSE), or Normalised Root Mean Squared Error (NRMSE).
@@
$RMSE(\mathbf{\hat{v}},\mathbf{v})$ $\triangleq\sqrt{MSE(\mathbf{\hat{v}},\mathbf{v})}$ (24)
$NMSE(\mathbf{\hat{v}},\mathbf{v})$ $\triangleq\frac{MSE(\mathbf{\hat{v}},\mathbf{v})}{MSE(\mathbf{\hat{v}},\langle\mathbf{\hat{v}}\rangle)}$ (25)
$NRMSE(\mathbf{\hat{v}},\mathbf{v})$ $\triangleq\sqrt{NMSE(\mathbf{\hat{v}},\mathbf{v})}$ (26)
A smaller error indicates better performance. Normalisation makes the NMSE and NRMSE measures dimensionless and independent of any scaling or units of the outputs, so are more comparable across experiments with different systems. Note that here normalisation is performed with respect to the target output, not the observed output (Lukoševičius,, 2012, p.661). This makes baselining a given system’s error using a constant observed value well-defined. In particular, NMSE = 1 and NRMSE = 1 if each $v_{t}$ is set to the mean of the target values $\langle\hat{\mathbf{v}}\rangle$. Experimental values greater than one indicate very poor perfomance; values less than one may also be achievable with just simple approaches (section 9.4.1).
Depending on what a given experiment is intended to convey, different choices of measure may be appropriate. To see how the difference between the measures, their behaviours are listed in table 8. This shows how normalising the error removes the effect of changing the standard deviation, which may be due to the choice of units or scaling.
std dev 0.1 1 2
MSE 0.009 0.813 3.713
RMSE 0.098 0.901 1.927
NMSE 1.007 1.002 1.039
NRMSE 1.003 1.001 1.019
@@ 9.3.2 Classification evaluation measures
For classification into a larger number of categories, a confusion matrix plot of predicted versus actual category shows both the proportion of correct predictions, and the distribution of incorrect predictions.
@@ 9.4 Presenting the results
9.4.1 Base cases
Calculate baseline success measures for your experiments.
@@ 9.4.1 Base cases
In time series based tasks, for example, a naive prediction is to use the previous target value, by setting $v(t)=\hat{v}({t-1})$. [^10] This is the persistence model of weather forecasting: tomorrow will be like today (AMS,, 2012). To illustrate this, we calculate the NRMSE of this base case for three popular benchmark tasks (table 9). This demonstrates that the baseline for success should be NMRSE $\lesssim 0.8$ for NARMA-10, and $\lesssim 0.4$ for Zurich sunspots.
@@ 10 Potential pitfalls of consistency
Different benchmarks require different behaviour from a Reservoir Computer (Dale et al.,, 2019). One can chose benchmarks that span this behaviour space, as well as benchmarks that have different approaches to data and what should be done with it.
@@ 11 Conclusion
Benchmarks are a useful way of evaluating Reservoir Computing in different contexts, particularly in terms of performance, or perhaps “usefulness”. Many of these benchmarks are inherited from the wider Machine Learning community, while others have been developed specifically for the field of Reservoir Computing. This provides a varied pool to choose from, and allows authors to choose benchmarks based on the argument they wish to present, while still placing their use of benchmarks in context with other works in the literature.
However, this mixture of sources has also led to a less well-defined benchmarking culture than in many other fields: there is no one “Reservoir Computing Benchmark Suite”, and even the closest equivalent, the NARMA benchmarks, has half a dozen different parameter sets and experimental setups that are inconsistent across different publications. This makes direct comparison between results of different works within the literature difficult, if not impossible.
New approaches such as CHARC (Dale et al.,, 2019) would allow us to survey a given substrate or reservoir model’s behaviour space, telling us whether a given substance would make a viable reservoir. While this is valuable work, however, with no direct link to benchmark tasks, it remains merely abstract. There is currently no direct mapping from benchmark tasks to areas within the behaviour space, and it is unclear if such a mapping would be possible.
@@
Acknowledgements
CW acknowledges support of a PhD studentship from the Department of Computer Science at the University of York. MT and SS acknowledge partial funding from the MARCH project, EPSRC grant number EP/V006029/1.

### v2

@@ 9.3.1 Time-series evaluation measures
With a few exceptions, dynamical systems-based benchmarks typically output a scalar value, rather than a multidimensional vector value, each timestep. Evaluation is then typically performed using one of Mean Squared Error (MSE), Root Mean Squared Error (RMSE), Normalised Mean Squared Error (NMSE), or Normalised Root Mean Squared Error (NRMSE). However, other measures may also be used, such as Mean Absolute Error (MAE) and Mean Absolute Percentage Error (MAPE).
@@
[…]
A smaller error indicates better performance. Normalisation makes the NMSE, NRMSE and MAPE measures dimensionless and independent of any scaling or units of the outputs, so are more comparable across experiments with different systems.
Note that here normalisation is performed with respect to the target output, not the observed output [22, p.661]. This makes baselining a given system’s error using a constant observed value well-defined. In particular, the normalised measures NMSE, NRMSE, and MAPE are equal to 1 if each $v_{t}$ is set to the mean of the target values $\langle\hat{\mathbf{v}}\rangle$. Experimental values greater than one indicate very poor performance; values less than one may also be achievable with just simple approaches (section 9.4).
Depending on what a given experiment is intended to convey, different choices of measure may be appropriate. To see how the difference between the measures, their behaviours are listed in table 8. This shows how normalising the error removes the effect of changing the standard deviation, which itself varies with the choice of units or scaling.
std dev 0.1 0.5 1 2
MSE 0.010 0.269 0.910 3.610
RMSE 0.101 0.519 0.954 1.900
NMSE 1.004 1.006 1.004 1.045
NRMSE 1.002 1.003 1.002 1.022
MAE 0.081 0.440 0.745 1.533
MAPE 100% 100% 100% 100%
@@ 9.3.2 Classification evaluation measures
For classification into a larger number of categories, symbol error rate (the fraction of incorrect symbols obtained) [25] may be used. A confusion matrix plot of predicted versus actual category shows both the proportion of correct predictions, and the distribution of incorrect predictions.
@@ 9.4 Presenting the results
Calculate baseline success measures.
@@
In time series based tasks, for example, a naive prediction is to use the previous target value, by setting $v(t)=\hat{v}({t-1})$. [^10] This is the persistence model of weather forecasting: tomorrow will be like today [134]. To illustrate this, we calculate the NRMSE of this base case for three popular benchmark tasks (table 9). This demonstrates that the baseline for success should be NMRSE $\lesssim 0.8$ for NARMA-10, and $\lesssim 0.4$ for Zurich sunspots.
@@ 10 Potential pitfalls of consistency
Different benchmarks require different behaviour from a Reservoir Computer [5]. One can choose benchmarks that span this behaviour space, as well as benchmarks that have different approaches to data and what should be done with it.
@@ 11 Conclusion
Benchmarks are a useful way of evaluating Reservoir Computing in different contexts, particularly in terms of performance, or perhaps ‘usefulness’. Many of these benchmarks are inherited from the wider Machine Learning community, while others have been developed specifically for the field of Reservoir Computing. This provides a varied pool to choose from, and allows authors to choose benchmarks based on the argument they wish to present, while still placing their use of benchmarks in context with other works in the literature.
However, this mixture of sources has also led to a less well-defined benchmarking culture than in many other fields: there is no one ‘Reservoir Computing Benchmark Suite’, and even the closest equivalent, the NARMA benchmarks, has half a dozen different parameter sets and experimental setups that are inconsistent across different publications. This makes direct comparison between results of different works within the literature difficult, if not impossible.
New approaches such as CHARC [5] would allow us to survey a given substrate or reservoir model’s behaviour space, telling us whether a given substance would make a viable reservoir. While this is valuable work, however, with no direct link to benchmark tasks, it remains merely abstract. There is currently no direct mapping from benchmark tasks to areas within the behaviour space, and it is unclear if such a mapping would be possible.
@@
Declarations
Ethical Approval
Not applicable: no human or animal research involved.
Competing Interests
The authors report there are no competing interests to declare.
Funding
C.W. acknowledges support of a PhD studentship from the Department of Computer Science at the University of York. M.T. and S.S. acknowledge partial funding from the MARCH project, EPSRC grant number EP/V006029/1.